The diagonal of a square is 3 ft. long, what is its perimeter?
Answer by stanbon(75887)

The diagonal of a square is 3 ft. long, what is its perimeter?
---------------------
Let each side of the square be "x"::
Equation:
2x^2 = 3^2
x^2 = (1/2)3^2
---
x = 3/sqrt(2)
------
Perimeter = 4x = 4(3/sqrt(2)) = 12sqrt(2)/2 = 6sqrt(2)
-----------------
Cheers,
Stan H.
==============